Summary information from the financial statements of two companies competing in the same industry follows.
  

Barco
Company
Kyan
Company
Barco
Company
Kyan
Company
Data from the current year-end balance sheets Data from the current year’s income statement
Assets Sales $ 770,000 $ 880,200
Cash $ 19,500 $ 34,000 Cost of goods sold 585,100 632,500
Accounts receivable, net 46,500 64,600 Interest expense 7,900 13,000
Merchandise inventory 84,440 132,500 Income tax expense 14,800 24,300
Prepaid expenses 5,000 6,950 Net income 162,200 210,400
Plant assets, net 290,000 304,400 Basic earnings per share 4.51 5.11
Total assets $ 445,440 $ 542,450 Cash dividends per share 3.81 3.93
Liabilities and Equity Beginning-of-year balance sheet data
Current liabilities $ 61,340 $ 93,300 Accounts receivable, net $ 29,800 $ 54,200
Long-term notes payable 80,800 101,000 Merchandise inventory 55,600 107,400
Common stock, $5 par value 180,000 206,000 Total assets 398,000 382,500
Retained earnings 123,300 142,150 Common stock, $5 par value 180,000 206,000
Total liabilities and equity $ 445,440 $ 542,450 Retained earnings 98,300 93,600

Required:
1a. For both companies compute the (a) current ratio, (b) acid-test ratio, (c) accounts receivable turnover, (d) inventory turnover, (e) days' sales in inventory, and (f) days' sales uncollected. (Do not round intermediate calculations.)
1b. Identify the company you consider to be the better short-term credit risk.
Answer #1
a)     Current ratio
Company Choose Numerator / Choose Denominator = Current ratio
current assets / Current liabilities = Current ratio
Barco 155440 / 61,340 = 2.5
Kyan 238050 / 93,300 = 2.6
b) Acid test ratio
Company Choose Numerator / Choose Denominator = Current ratio
Cash + short term investment + Current receivable / Current liabilities = Acid test ratio
Barco 19,500 46,500 / 61,340 = 1.1
Kyan 34,000 64,600 / 93,300 = 1.1
C) Accounts Receivable Turnover
Company Choose Numerator / Choose Denominator = Current ratio
Net sales / Average account receivable,net = Accounts receivable turnover
Barco 770,000 / 38,150 = 20.2 times
Kyan 880200 / 59,400 = 14.8 times
d) Inventory Turnover
Company Choose Numerator / Choose Denominator = Current ratio
Cost of goods sold / Average inventory = inventory turnover
Barco 585,100 / 70,020 = 8.4 times
Kyan 632500 / 119,950 = 5.3 times
e) Days' sale in inventory
Company Choose numerator / Choose Denominator * Days = Days' sales in inventory
Merchandise inventory / Cost of goods sold * 365 = Days' sales in inventory
Barco 84,440 / 585,100 365 = 52.7 days
Kyan 132,500 / 632500 365 = 76.5 days
e) Days' sale Uncollected
Company Choose numerator / Choose Denominator * Days = Days' sales in inventory
Current receivables,net / Net sales * 365 = Days' sale uncollected
Barco 46,500 / 770,000 365 = 22.0 days
Kyan 64,600 / 880200 365 = 26.8 days
Better short-term credit risk Barco